how much dextrose 70% should be mixed with 500 mL of dextrose 10% to make a 45% dextrose solution??
Im coming up with 272 Ml of the 70%. My book says it should be 700 mL..Typo??

Answer by stanbon(75887)

how much dextrose 70% should be mixed with 500 mL of dextrose 10% to make a 45% dextrose solution??
Im coming up with 272 Ml of the 70%. My book says it should be 700 mL..Typo??
-------------------------
Equation:
dex + dex = dex
-----
0.7x + 0.1*500 = 0.45(x+500)
------
70x + 10*500 = 45x + 45*500
-----------------------------
25x = 35*500
x = (7/5)(500)
x = 700 ml (amt. of 70% solution needed)
================================
Cheers,
Stan H.
---------------------